OpenStudy (anonymous):

A die is tossed. Find the odds against rolling a number greater than 4. *the probability of rolling a # greater than 4 is 1/3. *the probability that the event doesn't occur is 2/3. What is the ratio?

OpenStudy (anonymous):

there are six numbers in the die 1-6, right? and, how many numbers are there that is greater than 4 and is in die? what do you think ratio is now?

OpenStudy (anonymous):

2/6. Do I reduce it to 1/3? And would the ratio be 3:1?

OpenStudy (anonymous):

2/6 is 1/3 when you reduce, so there is you r answer.

OpenStudy (anonymous):

so the ratios is 1:3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what about the probability that the event doesn't occur, which is 2/3?

OpenStudy (anonymous):

probabiliy that the event doesn't occur is 1 - probabily that occurs. thus 1-1/3 = 2/3

OpenStudy (anonymous):

so the formula is 1/3 divided by 2/3 = 1/2. Right?

OpenStudy (anonymous):

what forumula?

OpenStudy (anonymous):

shasta you are correct

OpenStudy (anonymous):

and good job saruz